Last Saturday was another day of rewiring the controls for Tower 2. The wiring seemed to be a straight forward project, pharm but the process of running wires one at a time left other other original wiring bare and subject to a short. All the wiring will need to be replaced one at time, case for the secondary coleostat also.
The secondary coleostat is not a crucial to run electrically, as it effects solar declination only, and can be easily adjusted manually. Nevertheless, all deficient wiring needs to be replaced. There are some questions as to the wiring of some controls for the secondary coleostat, and most probably a wiring diagram after completion would be helpful.
Once the weather improves, the telescope will be put into use, and the the problems of focus will be addressed. The culminating lenses need to be cleaned. Negotiations with Richardson Gratings continue, and Debra has started a program to look at funding the new grating.
One of the larger projects will be to switch the primary and secondary mirrors. We have current polished ones, and once we are getting images, this will become more crucial. We will need a lot of muscle that day, so planning and attendance will be critical.